OpenStudy (anonymous):

What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12?

OpenStudy (anonymous):

for an arithmetic sequence you can write an explicit equation a(n) = a(1) - 6n so we can use the fact that a(11) = 12 to find what a(1) is in our sequence 12 = a(1) - 6(11) 12 = a(1) - 66 78 = a(1) now we have the equation a(n) = 78 - 6n from here if you want the 43rd term just use n=43